OpenStudy (anonymous):

Simplify..... X 4/5 over X 1/3

OpenStudy (anonymous):

X overX can cancel each other out [(4/5) / (1/3)=(4/5)*(3/1)=(4*3) / (5*1)=(12) / (5) \approx 2.4\]